Cost of Units Completed and in Process

The charges to Work in Process—Assembly Department for a period, together with information concerning production, are as follows. All direct materials are placed in process at the beginning of production.

Work in Process—Assembly Department
Bal., 3,000 units, 45% completed 7,800 To Finished Goods, 69,000 units ?
Direct materials, 71,000 units @ $1.7 120,700
Direct labor 103,800
Factory overhead 40,365
Bal., ? units, 20% completed ?

Cost per equivalent units of $1.70 for Direct Materials and $2.10 for Conversion Costs.

a. Based on the above data, determine the different costs listed below.

If required, round your interim calculations to two decimal places.

1. Cost of beginning work in process inventory completed this period $
2. Cost of units transferred to finished goods during the period $
3. Cost of ending work in process inventory $
4. Cost per unit of the completed beginning work in process inventory (Rounded to the nearest cent.) $

b. Did the production costs change from the preceding period?
Yes

c. Assuming that the direct materials cost per unit did not change from the preceding period, did the conversion costs per equivalent unit increase, decrease, or remain the same for the current period?
Increase
Answer #1

a1) Cost of Beginning WIP Inventory Completes this Period :-

= Beginnig WIP + Conversion Cost Incurred During Period

= $7800 + ($2.1*1650)

= $11,265

Conversion Equialent Unit :-

= 3000*55%

= 1650 Units

a2) Cost of units transferred to finished goods during the period :-

= Units Started and Commpleted Cost during the Period + Cost of Beginning WIP Inventory

= ((66000*$1.7)+(66000*$.2.1)) + $11,265

= $250,800 + $11,265

= 262,065

Unit Transferred to Finished Goods = Finished Goods - Beginning Balance of Inventory

= 69000 - 3000

= 66000 units

a3) Ending Unit = 3000+71000-69000 = 5000 units

Cost of Ending WIP Inventory = Material + Conversion

= (5000*$1.7) + (5000*$2.1)

= $8500 + $10500

= $19,000

a4) Cost Per Unit of the completed beginning WIP :-

= Cost of Beginning WIP/ Units of beginning WIP Inventory

= $11,265 / 3000

= $3.75 per unit

b) Yes, Production Cost is Changed.

Production Cost = $1.70 + $2.10 = $3.8 Per unit

c) Previous period conversion Cost = $3.75 - $1.7

= $2.05

Conversion Cost per Equivalent Unit Increase for the current period.